OpenStudy (anonymous):

The height, s, of a ball thrown straight down with initial speed 64 ft/sec from a cliff 80 feet high is s(t) = –16t2 – 64t + 80, where t is the time elapsed that the ball is in the air. What is the instantaneous velocity of the ball when it hits the ground? 256 ft/sec –96 ft/sec 0 ft/sec 112 ft/sec

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Figure out the time it hits the ground, then plug that into the derivative.

OpenStudy (anonymous):

It hits the ground when \(s(t)=0\).

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@arnavguddu

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

have u studied derivative?

OpenStudy (anonymous):

@familyguymath

OpenStudy (anonymous):

yas

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

well then, \[v(t) = \frac{ d }{ dx } s(t) = -32t - 64 \] \[a(t) = \frac{ d }{ dx } v(t) = -32 \] and use this formula: \[v^2 - v_{0}^2 = 2a \Delta x\]

OpenStudy (anonymous):